OpenStudy (anonymous):

Differentiate the following:

OpenStudy (anonymous):

\[\cos(x ^{2}+2y)=xe ^{y ^{2}}\]

OpenStudy (sidsiddhartha):

use chain rule

OpenStudy (anonymous):

For both parties?

OpenStudy (sidsiddhartha):

yeah

OpenStudy (sidsiddhartha):

in right hand side product rule

mathslover (mathslover):

for both "parties" .. lol!

OpenStudy (anonymous):

\[\cos(2x+2) + \sin(x ^{2}+2y) = xe ^{y ^{2}}(2y)+e ^{y ^{2}}\] right?

OpenStudy (sidsiddhartha):

\[\frac{ d }{ dx}\cos(x^2+2y)=[-\sin(x^2+2y)*\frac{ d }{dx }(x^2+2y)]\]

OpenStudy (sidsiddhartha):

so \[LHS=[-\sin(x^2+2y)*(2x+2\frac{ dy }{ dx })]\] ok with this

OpenStudy (anonymous):

But I think it will be a plus in product rule not Multiplication \[\frac{ d }{ dx}\cos(x^2+2y)=[-\sin(x^2+2y)+\frac{ d }{dx }(x^2+2y)]\]

OpenStudy (sidsiddhartha):

no the chain rule states that \[\large \large \frac{ d }{ dx }[f[g(x)]=f ^{'}[g(x)]*g^{'}(x)\]

OpenStudy (sidsiddhartha):

yes for product rule there will be plus but for chain rule there will be multiplication

OpenStudy (anonymous):

aha .. sorry about that so we have finished LHS ?

OpenStudy (sidsiddhartha):

ok LHS is finished now for RHS you made a bit mistake it will be \[\frac{ d }{ dx }[x*e^{y^{2}}]=x*\frac{ d }{ dx }[e^{y^{2}}]+e^{y^{2}}*\frac{ d }{ dx }(x)\] ok??

OpenStudy (sidsiddhartha):

now \[Rhs=x*[e^{y^{2}}*2y*\frac{ dy }{ dx}]+e^{y^{2}}\] ok with this??

OpenStudy (sidsiddhartha):

u forgot to write dy/dx ,because we are differentiating with respect to x

OpenStudy (anonymous):

product rule right?

OpenStudy (sidsiddhartha):

yeah and when ure differentiating e^(y^2) u have to use chain rule again \[\frac{ d }{ dx }[e^{y^{2}}]=e^{y^{2}}*\frac{ d }{ dx }(y^2)=e^{y^{2}}*2y \frac{ dy }{ dx }\]

OpenStudy (anonymous):

yeah I know that .. and what's next ?

OpenStudy (sidsiddhartha):

now both sides are completed now take the terms with (dy/dx) in one side and other terms in one side then take (dy/dx) common and u'll be finished

OpenStudy (anonymous):

ok wait me I will do and I will show you ..

OpenStudy (anonymous):

\[2x+2y \prime . 2y.y \prime = \sin(x ^{2}+2y).x.e ^{y ^{2}}+e ^{y ^{2}}\] so will come like that?

OpenStudy (sidsiddhartha):

ur doing some thing wrong check again

OpenStudy (anonymous):

\[2y \prime . 2y.y \prime = \sin(x ^{2}+2y).x.e ^{y ^{2}}+e ^{y ^{2}}-2x\] now it's true?